OpenStudy (anonymous):

Determine the amount needed to be deposited today(as a one time deposit) in order to have $35,000 four years from now. At 8% Solve for: Compounded: Annually Semi Annually and Quarterly

OpenStudy (e.mccormick):

So a basic compounding without any added investments. Know the formula?

OpenStudy (e.mccormick):

OK, you see the formula and you have expressed you have a problem with "Annually Semi Annually and Quarterly" So the issue is translating what the question is asking, right?

OpenStudy (anonymous):

yes, when it comes to words/story problems I have a very difficult time. I looked at the formula you sent, but I don't understand it. I see things differently, I need an answer and then I tend to work the problem backwards, that is how I do math...strange I know.

OpenStudy (e.mccormick):

Actually, they gave you the answer this time. They want you to find P. \[A=P\left(1+\frac{r}{n} \right)^{nt}\] P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year That last one, n, is the key to those terms. Anniually means n=1 Semi Anniually means n=2 Quarterly means n=4 Now, "A = amount of money accumulated after n years, including interest." That is where you want to be at. They give that to you as $35,000.

OpenStudy (e.mccormick):

So, start replacing things in:\[A=P\left(1+\frac{r}{n} \right)^{nt}\]With what was given, and check with me if you have the right thing in the right place.

OpenStudy (anonymous):

A= 35,000(1 + 4/1)

OpenStudy (e.mccormick):

They are giving you A, they want you to solve for P. Also, a problem with r and t there. r is the rate, so the PERCENTAGE rate, but as a decimal. That help? And t is Time in Years.

OpenStudy (anonymous):

35,000*.08*4 ?

OpenStudy (e.mccormick):

Well, closer. I will use the math thingie here. \[35000=P\left(1+\frac{.08}{n} \right)^{n4}\] Now, the n depends on which one of the three you are doing. So 1 for the annually. You need to solve for P.

OpenStudy (anonymous):

I am completely lost...this is why it takes me so long to pass math, I am so bad with it. I have been here trying to figure it out and I have become frustrated.I am the type that has to see it done, and then I try doing it. Thank you for your time, I really appreciate it.

OpenStudy (e.mccormick):

Well, then let me look at the first one. We have this right now: \[35000=P\left(1+\frac{.08}{n} \right)^{n4}\] And n is 1 for the first one, so: \[35000=P\left(1+\frac{.08}{1} \right)^{1\times 4}\] Which simplifies pretty easy: \[35000=P\left(1+.08 \right)^{4}\] And a little more: \[35000=P\left(1.08 \right)^{4}\] Now, to solve for P, we need P alone, so I need to divide off what is multiplying P. \[\frac{35000}{\left(1.08 \right)^{4}}=P\] And that can go into some calculators, but if not, I take one more step, do the power. \[\frac{35000}{1.3605}=P\]Which is rouded off to enough decimals that it won't cause a problem.

OpenStudy (anonymous):

25,725 ?

OpenStudy (e.mccormick):

I got 25,726.04 for the first one. So basically the same.

OpenStudy (anonymous):

ok..so let me try semi

OpenStudy (e.mccormick):

kk.

OpenStudy (anonymous):

35000/1.8509=18909.7

OpenStudy (e.mccormick):

Hmmm... Remember, it only simplified so much because n was 1. When n is 2, it changes things.

OpenStudy (e.mccormick):